OpenStudy (anonymous):

What is the inverse of y = 3x + 9?

OpenStudy (anonymous):

y-9=3x =>x=(y-9)/3

OpenStudy (anonymous):

you just replace x by y and y by x to get the inverse function. and then solve for y . to your problem, when replace you have x= 3y + 9 ---> 3y =x-9---> y = (x -9)/3 or y = 1/3x -3 . that is inverse function of your function hope this help